OpenStudy (anonymous):

The surface area of this pyramid is one hundred eighty-nine square centimeters. The area of the square base is forty-nine square centimeters. What is the area of one triangular face of the lateral surface? Show the steps of your solution and explain your work.

OpenStudy (anonymous):

@LaShaun97

OpenStudy (anonymous):

A square pyramid is formed by a square base and four triangular faces surface area of pyramid = area of square + 4(area of triangle) 189 = 49 + 4(area of triangle) area of triangle = (189 - 49) /4 = 140/4 = 35 square cms

OpenStudy (anonymous):

0.0 is that my answer?

OpenStudy (anonymous):

no i was just explaining it lol hold on

OpenStudy (anonymous):

ohhh haha I have my blonde moments

OpenStudy (anonymous):

lol i see

OpenStudy (anonymous):

The surface area of the pyramid is 189. Subtract from that the surface area of the base (49) and you get 140, which is the total surface area of the four triangular sides. Assuming that the base is horizontal, and the vertex (top) is directly above the middle of the base, then all four triangular sides are congruent, i.e. they each have the same area. So if four equal triangles have a total area of 140, then each triangle has an area of 140 / 4 = 35

OpenStudy (anonymous):

your answer is 35

OpenStudy (anonymous):

I have to show the steps

OpenStudy (anonymous):

im really not sure how to show the work hold on let me get dmezz to help us

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i gtg

OpenStudy (anonymous):

iight